# CodeBrain: Bridging Decoupled Tokenizer and Multi-Scale Architecture for EEG Foundation Model

### v1

@@ Abstract
Electroencephalography (EEG) provides real-time insights into brain activity and is widely used in neuroscience. However, variations in channel configurations, sequence lengths, and task objectives limit the transferability of traditional task-specific models. Although recent EEG foundation models (EFMs) aim to learn generalizable representations, they struggle with limited heterogeneous representation capacity and inefficiency in capturing multi-scale brain dependencies. To address these challenges, we propose CodeBrain, an efficient EFM structurally aligned with brain organization, trained in two stages. (1) We introduce a TFDual-Tokenizer that independently tokenizes heterogeneous temporal and frequency components, enabling a quadratic expansion of the discrete representation space. This also offers a degree of interpretability through cross-domain token analysis. (2) We propose the EEGSSM, which combines a structured global convolution architecture and a sliding window attention mechanism to jointly model sparse long-range and local dependencies. Unlike fully connected Transformer models, EEGSSM better reflects the brain’s small-world topology and efficiently captures EEG’s inherent multi-scale structure. EEGSSM is trained with a masked self-supervised learning objective to predict token indices obtained in TFDual-Tokenizer. Comprehensive experiments on 10 public EEG datasets demonstrate the generalizability of CodeBrain with linear probing. By offering biologically informed and interpretable EEG modeling, CodeBrain lays the foundation for future neuroscience research. Both code and pretraining weights will be released in the future version.
@@ 1 Introduction
Electroencephalography (EEG) data represent the brain’s electrical activity, captured by placing electrodes on the scalp niedermeyer2005electroencephalography . With its high temporal resolution, EEG provides real-time insights into brain dynamics, supporting neuroscience and cognitive research da2013eeg ; binnie2003cognitive . To enable automated analysis, researchers have developed various task-specific models, including sleep staging phan2021xsleepnet ; lee2025explainable ; ma2024st , emotion recognition jia2020sst ; guo2024eeg ; wang2024dmmr , motor imagery wang2024st ; ari2024nf ; li2020learning , and other applications guerra2024exploring ; hu2024exploring . However, these models are usually trained from scratch and tend to capture task-specific patterns, limiting their applicability beyond the original task. Moreover, variations in channel configurations and input lengths across EEG tasks further hinder knowledge transfer. To tackle these issues, EEG foundation models (EFMs) are developed to learn universal representations for diverse downstream tasks zhou2025brain .
@@
Inspired by masked self-supervised pretraining in natural language processing van2017neural ; devlin2019bert , current EFMs commonly adopt patch-wise representation learning: EEG signals are divided into fixed-length patches, encoded into latent representations, and trained to reconstruct the masked portions. This form of pretraining offers flexibility in downstream applications with varying channel configurations or input lengths by adjusting patch arrangement and number. Among reconstruction-based methods, some directly predict the masked raw signals wang2024eegpt ; wang2024cbramod , a task that is challenging due to EEG’s inherent noise and variability. Others adopt vector quantization to abstract away signal-level fluctuations jianglarge ; pradeepkumar2025single , and follow a two-stage pipeline: 1) Learn a tokenizer with a codebook to quantize EEG patches into discrete tokens, and 2) pretrain the EFM to predict the token indices of masked patches. By acting as a noise-robust latent space, the codebook enables the model to learn more expressive representations. Despite these advances, existing EFMs still face challenges.
@@
Failing to Decouple Heterogeneous EEG Representations. Recent EFMs adopt vector quantization with a single codebook for noise-robust representation jianglarge ; pradeepkumar2025single , following image-based VQ-VAE architecture van2017neural , where spatially entangled and homogeneous visual features make a single codebook sufficient zheng2023online ; mentzerfinite . However, EEG signals exhibit heterogeneous structures: temporal and frequency components reflect distinct aspects of brain activity miwakeichi2004decomposing . Applying a unified codebook to such heterogeneous data can blur domain-specific characteristics, potentially conflating localized temporal events with global frequency patterns liu2024vector and limiting the representation capacity.
@@
Struggling with Modeling Global Brain Dependencies. EEG signals exhibit sparse global dependencies and strong local correlations, reflecting the brain’s small-world topology bullmore2009complex ; bassett2006small ; he2009uncovering . Efficient modeling of such structure requires capturing relationships in a scalable way. However, most EFMs yang2023biot ; jianglarge ; mohammadi2024eeg2rep ; wang2024eegpt ; wang2024cbramod adopt Transformer architectures vaswani2017attention with fully connected self-attention. This over-connected design is misaligned with the brain’s sparse structure and struggles to efficiently capture global dependencies due to its quadratic complexity with sequence length tay2021long ; hong2025eegm2 ; tegon2025femba .
@@
Neglecting Local Dependencies within EEG Patches. EEG signals exhibit rich local waveform structures over short temporal windows, reflecting crucial transient neural events tatum2021handbook . However, most existing EFMs represent each EEG patch as a single token and apply attention mechanisms only at the patch level wang2024cbramod ; jianglarge , thereby ignoring important local dependencies within patches.
Figure 1: Overview of our contributions. (a) Decoupled vector quantization independently tokenizes temporal and frequency components to preserve heterogeneous EEG structures and enhance representation capacity. (b) State Space Model efficiently captures sparse global dependencies across patches. (c) Sliding Window Attention models fine-grained local dependencies within patches.
@@
To address the above challenges, we propose CodeBrain, an efficient EEG foundation model that integrates a decoupled time–frequency tokenizer and a multi-scale architecture. CodeBrain is trained in two stages. In the first stage, we introduce TFDual-Tokenizer (Fig. 1(a)), which independently tokenizes heterogeneous temporal and frequency EEG components. To stabilize training, we apply a contrastive learning objective to optimize the temporal codebook. In the second stage, we propose EEGSSM, a framework for learning multi-scale EEG representations by masked self-supervised learning. EEGSSM predicts token indices obtained from the TFDual-Tokenizer for masked patches. To address the inefficiency of over-connected attention and better align with the brain’s small-world structure, we introduce Structured Global Convolution (SGConv) (Fig. 1(b)) as the backbone. Inspired by Structured State Space Models gu2021efficiently ; smith2023simplified ; li2022makes ; gu2022parameterization , SGConv performs efficient global convolution over long EEG sequences, naturally capturing long-range dependencies and local clustering patterns characteristic of brain networks. Complementarily, we integrate a Sliding Window Attention (SWA) mechanism (Fig. 1(c)), adapted from long-sequence modeling beltagy2020longformer , to capture local intra-patch dependencies. Together, SGConv and SWA enable EEGSSM to perform efficient and scalable multi-scale EEG representation learning. Our detailed contributions can be summarized as follows:
  • Decoupled Heterogeneous Tokenizer. We propose the TFDual-Tokenizer, which independently encodes temporal and frequency components of EEG signals, leading to a quadratic expansion of the discrete representation space. It also provides interpretability, supported by visualization analyses in Section 4.5. To the best of our knowledge, this is the first vector quantization framework that decouples heterogeneous domains in EEG foundation models.
  • Efficient Multi-Scale Architecture. We design an efficient architecture combining a structured global convolution architecture backbone for global inter-patch dependencies and sliding window attention mechanism for local intra-patch dependencies. This design aligns with the brain’s small-world topology, enabling scalable multi-scale representation learning.
  • Comprehensive Validation. We evaluate CodeBrain on 8 diverse downstream brain-computer interface tasks using 10 public EEG datasets. Our CodeBrain model demonstrates strong generalization and state-of-the-art performance with linear probing.
@@ 2 Related Work
EEG Foundation Models. Inspired by foundation models in vision and language wang2023internimage ; achiam2023gpt , EEG research is shifting from task-specific models jia2020graphsleepnet ; wang2024caresleepnet ; chen2025long to EFMs for learning expressive representations. Current EFMs can be divided into two categories. 1) Contrastive learning-based (CL) chen2025large : BENDER kostas2021bendr first showed the ability of CL for EEG representations. Then, the Brant series zhang2023brant ; yuan2024brant ; zhang2024brant enables joint representation learning across physiological signals using CL. 2) Reconstruction-based: BIOT yang2023biot pioneers cross-modal pretraining for biosignals, including EEG. Subsequent models focus specifically on EEG, learning representations by predicting masked discrete tokens jianglarge ; jiangneurolm or reconstructing raw signals wang2024eegpt ; mohammadi2024eeg2rep ; wang2024cbramod . Most existing EFMs adopt Transformer architecture, which is suboptimal for EEG due to poor handling of sparse dependencies, quadratic complexity, and ignoring local dependencies by treating each patch as a token.
@@
EEG Tokenization. Tokenization has been key in NLP for generating generalizable and interpretable input representations sennrich2016neural ; kudo2018sentencepiece . Inspired by this, early EFMs used patch-based continuous tokenization yang2023biot ; yuan2024brant to handle EEG noise and variability, but without quantization, leading to unbounded and less interpretable representations. LaBraM jianglarge introduced vector quantization to learn discrete EEG tokens, following the VQ-VAE design from vision tasks van2017neural . However, this direct transfer overlooks EEG’s heterogeneous structure, limiting representation capacity. Moreover, the tokenizer of LaBraM is trained only on frequency-domain reconstruction due to convergence issues with raw signals reconstruction. Later efforts pradeepkumar2025single ; jiangneurolm all relied on a single codebook to encode mixed domain features, limiting both the representation space and interpretability.
@@
State Space Model Recent works have focused on enhancing classic State-Space Models (SSM) to more efficiently model sequential data using deep learning. For example, Syama et al. rangapuram2018deep used recurrent neural networks to learn parameters in SSM. However, the most significant progress came from Gu et al. gu2021efficiently with the introduction of the structured state space model (S4), which reduces the computational complexity of SSM in modeling long sequences using a special state transition matrix gu2020hippo . These low-rank and normal matrices enable SSM to compute global convolution kernels efficiently through fast Fourier transform across the entire sequence. Subsequently, some works have further improved the shortcomings of S4 in areas such as model architecture smith2022simplified and convolution raghu2023sequential , and have started applying it to tasks such as natural language processing dao2022hungry and time series analysis zhou2023deep . Recently, some works have also been applied to the EEG data, Tran et al. tran2024eeg leveraging SSMs to detect dementia. They extract temporal information from EEG signals through the Mamba architecture and combine it with frequency domain features to better manage the complexity of multivariate EEG. In the work of Gui et al. gui2024eegmambabidirectionalstatespace , SSM has also become the backbone network of the EEG foundation model. This further highlights the fast reasoning speed and efficient memory usage of the SSM model when processing EEG signals.
@@ 3 Methodology
3.1 Preliminaries
Decoupled Tokenization. Given a normalized EEG patch $X\in\mathbb{R}^{L}$, where $L$ is the patch length, our goal is to learn a tokenizer that maps $X$ to discrete tokens using two separate codebooks: one for the temporal domain and one for the frequency domain. Specifically, let $V^{(t)}\in\mathbb{R}^{K\times D}$ and $V^{(f)}\in\mathbb{R}^{K\times D}$ denote the temporal and frequency codebooks, respectively, where $K$ is the vocabulary size and $D$ is the embedding dimension of each token. The tokenizer function is defined as: $z^{(t)},z^{(f)}=f_{\text{tokenizer}}(X)$, where $z^{(t)},z^{(f)}\in\{1,\dots,K\}$ are the indices of the selected codes from the temporal and frequency codebooks.
EEG Representation Learning. Given unlabeled EEG sequences $\mathcal{X}=\{X_{i}\}_{i=1}^{N}$, where each sequence $X_{i}\in\mathbb{R}^{C\times T\times N}$ consists of $C$ channels, a sampling rate $T$, and a duration of $N$ seconds, the goal is to train an encoder $f_{\text{enc}}:\mathbb{R}^{C\times T\times N}\rightarrow\mathbb{R}^{D}$ that produces latent representations $Z_{i}=f_{\text{enc}}(X_{i})$.
@@ 3.2 Model Architecture
We introduce CodeBrain, a two-stage pretraining framework designed to address the two problem statements described in Section 3.1. The model architecture of CodeBrain is illustrated in Fig. 2. In the first stage, we propose TFDual-Tokenizer to decouple temporal and frequency representations of EEG signals, converting the raw signals into noise-robust tokens with potential interpretability. In the second stage, we introduce the EEGSSM framework for EEG representation learning.
@@ 3.3 TFDual-Tokenizer Pretraining
Neural Encoder For each patched and normalized EEG input $X\in\mathbb{R}^{C\times T\times N}$, we apply the Discrete Fourier Transform (DFT) cooley1965algorithm to obtain its frequency representation:
$X[k]=\sum_{n=0}^{N-1}x_{n}\cdot e^{-j\frac{2\pi}{N}kn},\quad k=0,1,\dots,N-1,$ (1)
@@
where $X[k]$ denotes the $k$-th frequency component, and $n$ denotes the number of patches. $X[k]$ is then combined with the raw time-domain signal and fed into the TFConv module, where both are processed in parallel through stacks of convolutional, batch normalization, and ReLU layers. The outputs $e_{t}$ and frequency representation $X[K]$ are concatenated to form a time-frequency embedding $e_{p}=Concat\{e_{t},X[K]\}$. To obtain patch representation $\tilde{e}_{n}$, we then add a positional embedding $e_{\text{pos}}$ and feed them into TF-Encoder, a Transformer encoder:
$\tilde{e}_{n}=TransformerEncoder(e_{p}+e_{pos}).$ (2)
@@
Dual Tokenizer We use two separate tokenizers with distinct codebooks for the temporal and frequency domains, denoted as $v^{(t)}_{j},v^{(f)}_{j}\in\mathbb{R}^{D}$, where $j=1,\dots,K$. Given the patch representation $\tilde{e}_{i}$ from the neural encoder, each tokenizer independently selects the nearest code from its codebook:
$\hat{z}^{(t)}_{i}=\arg\min_{j}\|\tilde{e}_{i}-v^{(t)}_{j}\|^{2},\quad\hat{z}^{(f)}_{i}=\arg\min_{j}\|\tilde{e}_{i}-v^{(f)}_{j}\|^{2}.$ (3)
@@
Frequency Codebook Training To train the frequency codebook, we reconstruct two fundamental features—amplitude and phase—from the code embeddings. For each EEG patch, we apply the DFT (Eq. (1)) to obtain the complex frequency representation: $X[k]=\text{Re}\{X[k]\}+j\cdot\text{Im}\{X[k]\}$
where $\text{Re}\{X[k]\}$ and $\text{Im}\{X[k]\}$ are the real part and imaginary part respectively, then the amplitude and phase can be calculated as:
$A_{n}=\sqrt{\text{Re}(X[k])^{2}+\text{Im}(X[k])^{2}},\quad\phi_{n}=\arctan 2\left(\text{Im}(X[k]),\text{Re}(X[k])\right).$ (4)
@@
We use z-score normalization to ensure stable training. The code embedding $v^{(f)}_{\hat{z}_{n}}$, retrieved from the frequency codebook, is passed through the F-Decoder, which consists of a Transformer encoder followed by two linear layers:
$y_{n}^{A}=TF^{(f)}(MLP(\tilde{e}_{n})),\quad y_{n}^{P}=TF^{(f)}(MLP(\tilde{e}_{n})).$ (5)
@@
where $y_{n}^{A}$ and $y_{n}^{P}$ are the predicted amplitude and phase, respectively. $TF^{(f)}$ denotes a Transformer encoder. The frequency codebook’s training objective frequency codebook is the mean squared error (MSE) loss:
$\mathcal{L}_{n}^{f}=||y_{n}^{A}-A_{n}||^{2}_{2}+||y_{n}^{P}-\phi_{n}||^{2}_{2}.$ (6)
@@
Temporal Codebook Training Direct reconstruction of temporal features might lead to non-convergence jianglarge . To address this, we combine contrastive loss with reconstruction loss to train a temporal codebook. Inspired by studies on physiological signals kiyasseh2021clocs , we assume temporal dependencies exist between different EEG patches, especially within the same channel. For an EEG sample $X\in\mathbb{R}^{C\times T\times N}$, we split it into two halves of length $N/2$. We use TF-Encoder in Eq. (2) to obtain the representations of these two segments $X_{h1},X_{h2}\in\mathbb{R}^{C\times T\times N/2}$ separately:
$h_{n}=TransformerEncoder(X_{{n}}),X_{n}\in\mathbb{R}^{C\times(T\times N/2)}.$ (7)
@@
We encourage the latent representations of different segments within a single sample $h_{n}$ and $\tilde{h}_{n}$ to be similar, while making the latent representations of the same segment across different samples $h_{n}$ and $h_{k}$ as distinct as possible. We use SimCLR loss chen2020simple to train the temporal codebook:
$\mathcal{L}_{CL}=-\log\frac{\exp\left(\mathrm{sim}({h}_{n},\tilde{h}_{n})/\tau\right)}{\sum\limits_{k=1}^{2N}\mathbb{1}_{[k\neq i]}\exp\left(\mathrm{sim}({h}_{n},{h}_{k})/\tau\right)}\quad,$ (8)

### v2

# CodeBrain: Towards Decoupled Interpretability and Multi-Scale Architecture for EEG Foundation Model

@@ Abstract
Electroencephalography (EEG) provides real-time insights into brain activity and supports diverse applications in neuroscience. While EEG foundation models (EFMs) have emerged to address the scalability issues of task-specific models, current approaches still yield clinically uninterpretable and weakly discriminative representations, inefficiently capture global dependencies, and neglect important local neural events. We present CodeBrain, a two-stage EFM designed to fill this gap. In the first stage, we introduce the TFDual-Tokenizer, which decouples heterogeneous temporal and frequency EEG signals into discrete tokens, quadratically expanding the representation space to enhance discriminative power and offering domain-specific interpretability by suggesting potential links to neural events and spectral rhythms. In the second stage, we propose the multi-scale EEGSSM architecture, which combines structured global convolution with sliding window attention to efficiently capture both sparse long-range and local dependencies, reflecting the brain’s small-world topology. Pretrained on the largest public EEG corpus, CodeBrain achieves strong generalization across 8 downstream tasks and 10 datasets under distribution shifts, supported by comprehensive ablations, scaling-law analyses, and interpretability evaluations. Both code and pretraining weights will be released in the future version.
@@ 1 Introduction
Electroencephalography (EEG) captures brain activity via scalp electrodes (niedermeyer2005electroencephalography, ) and provides high temporal-resolution signals for neuroscience and cognitive research (da2013eeg, ). To enable automated analysis, researchers have developed various task-specific models for applications such as sleep staging (lee2025explainable, ; ma2025st, ), emotion recognition (jia2020sst, ; liu2024vbh, ), motor imagery (li2020learning, ; jia2020mmcnn, ), and other applications (guerra2024exploring, ; hu2024exploring, ). However, building separate models from scratch for each task is resource-intensive and limits scalability, as shared knowledge across tasks cannot be effectively leveraged. Moreover, variations in channel configurations and input lengths across EEG tasks further hinder knowledge transfer. To tackle these issues, EEG foundation models (EFMs) are developed to learn universal representations for diverse downstream tasks (zhou2025brain, ).
@@
Inspired by masked self-supervised pretraining in natural language processing (van2017neural, ; devlin2019bert, ), current EFMs commonly adopt patch-wise representation learning: EEG signals are divided into patches, encoded into latent representations, and trained to reconstruct the masked portions. While this approach offers flexibility across varying channel configurations and input lengths by adjusting the patch number and arrangement, direct raw-signal reconstruction (wang2024eegpt, ; wang2024cbramod, ) remains challenging due to the inherent noise and variability of EEG. To mitigate this, recent studies have introduced codebook-based tokenization (jianglarge, ; pradeepkumar2025single, ), which abstracts away low-level fluctuations and provides a more robust latent space. Despite these advances, existing EFMs still face critical limitations, calling for new architectures.
@@
Failing to Decouple Heterogeneous EEG for Domain-Specific Interpretability. Recent EFMs adopt vector quantization for noise-robust representation (jianglarge, ; pradeepkumar2025single, ), following the VQ-VAE framework originally designed for images, where homogeneous visual features make a single tokenizer sufficient (van2017neural, ; mentzerfinite, ). In contrast, EEG exhibits heterogeneous structures: temporal and frequency components reflect distinct aspects of brain activity (miwakeichi2004decomposing, ). As illustrated in Fig. 1(a), signals matched in one domain may diverge in the other. Therefore, a mixed tokenizer can conflate domain-specific patterns (liu2024vector, ), weakening representation capacity and producing tokens difficult to align with clinically interpretable neural events or spectral rhythms.
@@
Struggling with Efficiently Modeling Global Brain Dependencies. EEG signals exhibit sparse global dependencies and strong local correlations, reflecting the brain’s small-world topology (bullmore2009complex, ; bassett2006small, ; he2009uncovering, ). Efficient modeling of such structure requires capturing relationships in a scalable way. However, most EFMs (yang2023biot, ; jianglarge, ; wang2024eegpt, ; wang2024cbramod, ) adopt Transformer architectures with fully connected self-attention (vaswani2017attention, ). This over-connected design is misaligned with the brain’s sparse structure and struggles to efficiently capture global dependencies due to its quadratic complexity with sequence length (tay2021long, ; hong2025eegm2, ; tegon2025femba, ).
@@
Neglecting Local Dependencies within EEG Patches. EEG signals exhibit rich local waveform structures over short temporal windows, reflecting crucial transient neural events (e.g., sleep waveforms in Fig. 1(d)) (tatum2021handbook, ). However, most existing EFMs represent each EEG patch as a single token and apply attention mechanisms only at the patch level (wang2024cbramod, ; jianglarge, ), thereby ignoring important local dependencies within patches.
Figure 1: Rationale and overview of CodeBrain beyond existing EFMs. (a) EEG signals are heterogeneous, as patches matched in one domain may diverge in the other. (b) We then propose a decoupled tokenizer for domain-specific interpretable representations while expanding the representation space. (c-d) Inspired by the brain’s small-world topology, a multi-scale architecture further captures inter-patch dependencies efficiently, while modeling overlooked intra-patch neural events. (e) These designs deliver performance gains across 10 EEG datasets.
@@
To address the above challenges, we propose CodeBrain, a novel EEG foundation model that integrates a decoupled tokenizer for domain-specific interpretability with a brain-inspired multi-scale architecture. CodeBrain is trained in two stages. In the first stage, we introduce the TFDual-Tokenizer (Fig. 1(b)), which decouples temporal and frequency EEG components into discrete tokens to support domain-specific interpretability. In the second stage, we develop EEGSSM, a masked self-supervised framework inspired by the brain’s small-world topology. EEGSSM adopts a structured global convolution backbone, conceptually related to recent state-space sequence models (smith2023simplified, ; li2022makes, ; gu2022parameterization, ), for sparse and efficient global modeling with a sliding window attention (SWA) mechanism for capturing local neural events overlooked by prior studies(Fig. 1(c–d)). Our detailed contributions are summarized as follows:
  • Decoupled Tokenizer for Domain-Specific Interpretability. We propose the TFDual-Tokenizer, which decouples temporal and frequency EEG components into discrete tokens. This design quadratically expands the representation space, and qualitative analyses suggest that the resulting tokens align with neural events and spectral rhythms. A contrastive objective is further applied to the temporal branch to stabilize training. To the best of our knowledge, this is the first tokenizer in EFMs to provide domain-specific interpretability.
  • Brain Small-World Topology Inspired Multi-Scale Architecture. We design EEGSSM, a patch-wise self-supervised framework for EEG. Guided by the brain’s small-world topology, it employs structured global convolution to capture sparse long-range temporal dependencies and sliding window attention to model local neural events. In addition, dynamic positional embeddings are used to flexibly learn spatial channel correlations.
  • Strong Generalization and Comprehensive Validation. Pretrained on the largest publicly available EEG corpus, TUEG (obeid2016temple, ), CodeBrain achieves strong performance on 8 downstream tasks across 10 datasets (Fig. 1(e)) with distribution shifts in cohorts and channel configurations. This suggests the model design plays a central role in generalization. Comprehensive ablations, scaling-law analyses (3M–150M parameters), together with visualization and quantitative analyses, further confirm its robustness, scalability, and provide domain-specific interpretability.
@@ 2 Related Work
Inspired by foundation models in vision and language (wang2023internimage, ; achiam2023gpt, ), EEG research is shifting from task-specific models (jia2020graphsleepnet, ; wang2024caresleepnet, ; chen2025long, ) to EFMs for learning expressive representations. Current EFMs can be divided into two categories. 1) Contrastive learning-based (CL) (chen2025large, ): BENDER (kostas2021bendr, ) first showed the ability of CL for EEG representations. Then, the Brant series (zhang2023brant, ; yuan2024brant, ; zhang2024brant, ) enables joint representation learning across physiological signals using CL. 2) Reconstruction-based: BIOT (yang2023biot, ) pioneers cross-modal pretraining for biosignals, including EEG. Subsequent models focus specifically on EEG, learning representations by predicting masked discrete tokens (jianglarge, ; jiangneurolm, ) or reconstructing raw signals (wang2024eegpt, ; mohammadi2024eeg2rep, ; wang2024cbramod, ). Most existing EFMs adopt Transformer architecture, which is suboptimal for EEG due to poor handling of sparse dependencies, quadratic complexity, and ignoring local dependencies by treating each patch as a token.
EEG Tokenization.
@@
Tokenization has been key in NLP for generating generalizable and interpretable input representations (sennrich2016neural, ; kudo2018sentencepiece, ). Inspired by this, early EFMs used patch-based continuous tokenization (yang2023biot, ; yuan2024brant, ) to handle EEG noise and variability, but without quantization, leading to unbounded and less interpretable representations. LaBraM (jianglarge, ) introduced vector quantization to learn discrete EEG tokens, following the VQ-VAE design from vision tasks (van2017neural, ). However, this direct transfer overlooks EEG’s heterogeneous structure, limiting representation capacity. Moreover, the tokenizer of LaBraM is trained only on frequency-domain reconstruction due to convergence issues with raw signals reconstruction. Later efforts (pradeepkumar2025single, ; jiangneurolm, ) all relied on a single codebook to encode mixed domain features, limiting both the representation space and interpretability.
State Space Model
@@
Recent works have focused on enhancing classic State-Space Models (SSM) to more efficiently model sequential data using deep learning. For example, rangapuram2018deep used recurrent neural networks to learn parameters in SSM. However, the most significant progress came from gu2021efficiently with the introduction of the structured state space model (S4), which reduces the computational complexity of SSM in modeling long sequences using a special state transition matrix (gu2020hippo, ). These low-rank and normal matrices enable SSM to compute global convolution kernels efficiently through fast Fourier transform across the entire sequence. Subsequently, some works have further improved the shortcomings of S4 in areas such as model architecture (smith2022simplified, ) and convolution (raghu2023sequential, ), and have started applying it to tasks such as natural language processing (dao2022hungry, ) and time series analysis (zhou2023deep, ). Recently, some works have also been applied to the EEG data, Tran et al. (tran2024eeg, ) leveraging SSMs to detect dementia. They extract temporal information from EEG signals through the Mamba architecture and combine it with frequency domain features to better manage the complexity of multivariate EEG. In the work of gui2024eegmambabidirectionalstatespace , SSM has also become the backbone network of the EEG foundation model. This further highlights the fast reasoning speed and efficient memory usage of the SSM model when processing EEG signals.
@@ 3.1 Model Architecture
We introduce CodeBrain, a two-stage pretraining framework designed to learn interpretable and universal EEG representations (Fig. 2). The model is motivated by complementary goals: 1) domain-specific interpretability via decoupled tokenization of heterogeneous temporal and frequency information, achieved by the proposed TFDual-Tokenizer, and 2) multi-scale modeling of EEG sequences inspired by the brain’s small-world topology, addressed by the EEGSSM framework. This design lets Stage 1 learn a tokenizer of patch-level codes, while Stage 2 leverages it for EEG representations. We next provide a formal definition of the two stages to clarify their respective roles.
Stage 1: Decoupled Tokenization. Given a normalized EEG patch $\mathbf{x}\in\mathbb{R}^{L}$, where $L$ is the patch length, our goal is to discretize $\mathbf{x}$ into temporal and frequency tokens, enabling domain-specific representation learning. Specifically, let $Vt\in\mathbb{R}^{K\times D}$ and $Vf\in\mathbb{R}^{K\times D}$ denote the temporal and frequency codebooks, where $K$ is the vocabulary size and $D$ is the embedding dimension of each token. The tokenizer function is defined as: $vt,vf=f_{\text{tokenizer}}(\mathbf{x}),\quad vt,vf\in\mathbb{R}^{D}.$
Stage 2: EEG Representation Learning. Given unlabeled EEG sequences $\mathcal{X}=\{X_{m}\}_{m=1}^{N}$, where each $X_{m}\in\mathbb{R}^{C\times f\times T}$ consists of $C$ channels, sampling rate $f$, and $T$ seconds. We divide each sequence into $n$ non-overlapping patches of $t$ seconds, so each patch length is $L=f\cdot t$. The goal is to train an encoder $f_{\text{enc}}:\mathbb{R}^{C\times n\times L}\rightarrow\mathbb{R}^{C\times n\times D}$ that produces latent representations $Z_{m}=f_{\text{enc}}(X_{m})$.
@@ 3.2 TFDual-Tokenizer Pretraining
Neural Encoder For each patch $\mathbf{x}_{i}\in\mathbb{R}^{L}$, we apply the Discrete Fourier Transform (DFT) (cooley1965algorithm, ) to obtain its frequency representation:
$\mathbf{x}_{i}[k]=DFT(\mathbf{x}_{i})$ (1)
@@
where $\mathbf{x}_{i}[k]$ denotes the $k$-th frequency component. $\mathbf{x}_{i}[k]$ and the $\mathbf{x}_{i}$ are fed into the TFConv module, where they are processed in parallel through stacks of convolutional, batch normalization, and ReLU layers. The temporal representation $e^{t}_{i}=TFConv(\mathbf{x}_{i})$ and frequency representation $e_{i}^{f}=TFConv(\mathbf{x}_{i}[k])$ are concatenated to form a time-frequency embedding $e^{p}_{i}=Concat\{e^{t}_{i},e^{f}_{i}\}$.
To get patch representation $\tilde{e}_{i}$, we then add a positional embedding $e_{\text{pos}}$ and feed it into TF-Encoder:
$\tilde{e}_{i}=Encoder(e^{p}_{i}+e_{pos}).$ (2)
We choose a Transformer encoder here since this stage is patch-to-token, where its ability to model local contextual relations makes it well-suited for capturing patch-level patterns. To keep the tokenizer channel-agnostic, $e_{\text{pos}}$ is shared temporal embeddings without channel-specific identities.
@@
Dual Tokenizer We use two separate tokenizers with distinct codebooks for the temporal and frequency domains, denoted as $vt_{j},vf_{j}\in\mathbb{R}^{D}$, where $j=1,\dots,K$. Given the patch representation $\tilde{e}_{i}$ from the neural encoder, each tokenizer independently selects the nearest code from its codebook:
$pt_{i}=\arg\min_{j}\|\tilde{e}_{i}-vt_{j}\|^{2},\quad pf_{i}=\arg\min_{j}\|\tilde{e}_{i}-vf_{j}\|^{2},$ (3)
where $pt_{i}$ and $pf$ denote the closest positions for the embeddings in the temporal and frequency domain codebook. The effectiveness of the Dual Tokenizer is based on the following proposition:
Proposition 2.1 Decoupling temporal and frequency codebooks yields representations that are no less effective than those from a joint codebook.
Proof. See Appendix C.
For this Proposition, we provide empirical validation in Sections 4.4, 4.5 and analysis of Dual Tokenizer ’s interpretable structure in Appendix A.
@@
Frequency Codebook Training To train the frequency codebook, we reconstruct amplitude and phase from the code embeddings. For each EEG patch, we apply the DFT to obtain the frequency representation: $\mathbf{x}_{i}[k]=\text{Re}\{\mathbf{x}_{i}[k]\}+j\cdot\text{Im}\{\mathbf{x}_{i}[k]\}$ where $\text{Re}\{\mathbf{x}_{i}[k]\}$ and $\text{Im}\{\mathbf{x}_{i}[k]\}$ are the real part and imaginary part respectively, then the amplitude and phase can be calculated as:
$A_{i}=\sqrt{\text{Re}(\mathbf{x}_{i}[k])^{2}+\text{Im}(\mathbf{x}_{i}[k])^{2}},\quad\phi_{i}=\arctan 2\left(\text{Im}(\mathbf{x}_{i}[k]),\text{Re}(\mathbf{x}_{i}[k])\right).$ (4)
@@
We use z-score normalization to ensure stable training. The code embedding ${vf}_{i}$, retrieved from the frequency codebook, is passed through the F-Decoder, which consists of a Transformer encoder followed by two linear layers:
$y_{i}^{A}=Encoder(MLP(\tilde{e}_{i})),\quad y_{i}^{P}=Encoder(MLP(\tilde{e}_{i})).$ (5)
@@
where $y_{i}^{A}$ and $y_{i}^{P}$ are the predicted amplitude and phase, respectively. The frequency codebook’s training objective is the mean squared error (MSE) loss:
[…]
Temporal Codebook Training
@@
Direct reconstruction of temporal features might lead to non-convergence (jianglarge, ). To address this, we combine contrastive loss with reconstruction loss to train a temporal codebook. Inspired by studies on physiological signals (kiyasseh2021clocs, ), we assume temporal dependencies exist between EEG segments, especially within the same channel. For an EEG segment $X_{m}\in\mathbb{R}^{C\times n\times L}$, we split it into two halves of length $n/2$. We use TF-Encoder in Eq. (2) to obtain the representations of these two parts $X_{m1},X_{m2}\in\mathbb{R}^{C\times\frac{2}{n}\times L}$ separately:
$e^{h}_{mi}=Encoder(X_{mi}),i\in\{1,2\}.$ (7)
@@
We encourage the latent representations of different parts within a single segment $e^{h}_{m1}$ and $e^{h}_{m2}$ to be similar, while making those of the same part across different segments $e^{h}_{mi}$ and $e^{h}_{sk}$ as distinct as possible, where $X_{s}\in\mathbb{R}^{C\times n\times L},s\neq m,k\neq i$. SimCLR loss (chen2020simple, ) is used for training:
$\mathcal{L}^{CL}_{m}=-\log\frac{\exp\left(\mathrm{sim}(e^{h}_{m1},e^{h}_{m2})/\tau\right)}{\sum\limits_{k=1}^{2N}\mathbb{1}_{[k\neq i]}\exp\left(\mathrm{sim}(e^{h}_{mi},e^{h}_{sk})/\tau\right)},$ (8)